OpenStudy (anonymous):

compute the discriminant. then determine the number and types of solutions for the given equation. x^2-8x+7=0

OpenStudy (anonymous):

the discriminant > 0 ==> two different real solutions

OpenStudy (anonymous):

the discriminant =64-28 >0

OpenStudy (anonymous):

\[\sqrt{1+10\sqrt{x}}=1+\sqrt{x}\]